A health maintenance organization (HMO) wants to estimate the mean length of a hospital stay. How large a sample of patient records is necessary if the HMO wants to be 99 percent confident of the estimate and wants the estimate to be within plus or minus 0.2 days? An earlier study showed the standard deviation of the length of stay to be 0.25 days.
Answer #1
for99% CI crtiical Z          = 2.576 (from excel function : normsinv(0.995)
standard deviation σ= 0.25
margin of error E = 0.2
required sample size n=(zσ/E)2 =(2.576*0.25/0.2)^2     = 11